40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 151 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  9352   Wed Nov 6 08:33:53 2013 SteveUpdateIOOPoiting changes of PSL

Quote:

 Full list tomorrow: IP-Ang & Pos, ETMY-T, ETMY-Oplev, ETMX-T, IOO-Ang & Pos

 RA: No one in the control room this evening can understand what this ELOG means. Please use more words.

 Yesterday the last steering mirror mount on the PSL was changed, Manasa recovered the MC alignment and Jenne locked the arms.

 I centered the following qpds:  ASC-IBQPD, LSC-TRY, SUS-ETMY_OPLEV, LSC-TRX, SUS_ETMX_OPLEV

 Touching the PSL pointing IOO-QPD_ANG & POS was a mistake. We lost the reference of the well refined MC input.

 

One and 20 days TRENDS  plot showing the PSL output drift in pitch can be power drop

However initial pointing is amazingly good. ( I wonder about the lens in front of the qpd ?)

Attachment 1: 1dayTRENDofPOITING.png
1dayTRENDofPOITING.png
Attachment 2: 20dayTRENDioo.png
20dayTRENDioo.png
  9351   Tue Nov 5 19:55:12 2013 ranaUpdateSUSoplev XY-plots reflect new calibration

I used the same OSEM SUSPIT/YAW method as before to calibrate the ETMY optical lever signals. They were off by a factor of ~10.

ETMY Pitch     300  /  26    (old/new)     urad/counts

ETMY Yaw       300  /  31    (old/new)     urad/counts

These should be redone with the Kakeru / Ottaway arm cavity power technique if we want to get better than ~30% accuracy.

  9350   Tue Nov 5 19:50:07 2013 manasaUpdateElectronicsIOO rack +/-5V power supplies

The power supply to the ADC box on the IOO rack (that reads the beat I & Q signals) was pulled out because it did not run through any fuse and was connected directly to the power supply. 

There were already connections running from the +/-5 V power supply. They were powering the mode cleaner demod board rack. In order to remove the ADC power connectors from the power supply, I notified Jenne in the control room because turning off the power supply would affect the MC. I switched off the +/-5V power supplies at the same time. The ADC power connectors were removed. The +/-5V power supplies were then turned ON again at the same time. Jenne relocked the MC after this.

I have still not connected the ADC to the fuse rack power supply because this requires the +/-5V power supplies to be turned OFF again in order to pull out new connections from the fuse rack and I need to make a new ADC power connector with thicker wires.

  9349   Tue Nov 5 19:39:27 2013 JenneUpdateLSCOpLev time series

[Rana, Jenne]

We looked at the time series for all the oplevs except the BS, from last Tuesday night, during a time when we were building up the power in the arms.  We conclude from a 400 second stretch of data that there is not discernible difference in the amount of motion of any optic, when the cavities are at medium power, and when they're at low power.  Note however, that we don't have such a nice stretch of data for the really high powers, so the maximum arm power in these plots is around 5.  Both the TRX and TRY signals look fairly stationary up to powers of 1 or 2, but once you get to 4 or 5, the power fluctuations are much more significant.  So, since this isn't caused by any optic moving more, perhaps it's just that we're more sensitive to optic motion when we're closer to resonance in the arms.

However, from this plot, it looks like the ETMY is moving much more than any other optic.  On the other hand, ETMY has not ever been calibrated (there's an arbitrary 300 in there for the calibration numbers on the ETMY oplev screen).  So, perhaps it's not actually moving any more than other optics.  We should calibrate the ETM oplevs nicely, so we have some real numbers in there.  ETMX also only is roughly calibrated, relative to the OSEMs.  We should either do the move-the-QPD calibration, or a Kakeru-style pitch and yaw some mirrors and look at transmitted power.

Traces on this xml file have been filtered with DTT, using zpk([0],[0.03],1,"n").

OpLevs_during_PRMI_2arms.pdf

 

  9348   Tue Nov 5 17:12:48 2013 JenneFrogsLSCillegal power supply about to expire

Quote:

 Is this your illegally installed HP bench power supply?

 Steve has promised to add another row of fuses to the LSC rack first thing in the morning.  Then, during Wednesday Chores, we can move the wires from the power supply to the fused power.

STEVE:  NEVER MIND about doing this in the morning.  Let's chat at the lunch meeting about what needs to be done to power things down, then back up again, in a nice order, and we can do it after lunch.  

So, please do not do anything to the LSC rack tomorrow!  Thank you.

  9347   Tue Nov 5 17:12:34 2013 SteveUpdateLSCcentered qpds

 Full list tomorrow: IP-Ang & Pos, ETMY-T, ETMY-Oplev, ETMX-T, IOO-Ang & Pos

 RA: No one in the control room this evening can understand what this ELOG means. Please use more words.

Attachment 1: recentered.png
recentered.png
  9346   Tue Nov 5 16:47:19 2013 ranaFrogsLSCillegal power supply about to expire

 Is this your illegally installed HP bench power supply?

20131023_222351.jpg

If so, or if not but you care about the signal that passes through these amplifiers, I suggest you remove this temporary power supply and wire the power from the rack power supplies through the fuse blocks and possibly use a voltage regulator.

In 24 hours, that power supply will be disconnected and the wires snipped if they are still there.

  9345   Tue Nov 5 16:47:09 2013 JenneUpdateLSCEnd transmission QPDs

I think Steve is trying to align the end transmission QPDs, since the arms are locked nicely right now.  I noticed that the QPDX pitch and yaw signals were digital zeros.  A quick look determined that the QPDX matrix to go from 4 quadrants to 3 degrees of freedom had been filled in for the POS row, but not pitch and yaw.  So, I copied the QPDY matrix over to QPDX (so the ordering of the rows and columns is assumed to be the same). 

Hopefully this will get us close to centered, but I suppose we ought to check really which quadrant is which, by shining a laser pointer at each quad at each end.

  9344   Tue Nov 5 16:39:54 2013 GabrieleUpdateLSCThoughts on Transition to IR

Quote:

Gabriele and I talked for a while on Wednesday afternoon about ideas for transitioning to IR control, from ALS. 

I think one of the baseline ideas was to use the sqrt(transmission) as an error signal.  Gabriele pointed out to me that to have a linear signal, really what we need is sqrt( [max transmission] - [current transmission] ), and this requires good knowledge of the maximum transmission that we expect.  However, we can't really measure this max transmission, since we aren't yet able to hold the arms that close to resonance.  If we get this number wrong, the error signal close to the resonance won't be very good.

Gabriele suggested maybe using just the raw transmission signal.  When we're near the half-resonance point, the transmission gives us an approximately linear signal, although it becomes totally non-linear as we get close to resonance.  Using this technique, however, requires lowering the finesse of PRCL by putting in a medium-large MICH offset, so that the PRC is lossy.  This lowering of the PRC finesse prevents the coupled-cavity linewidth of the arm to get too tiny.  Apparently this trick was very handy for Virgo when locking the PRFPMI, but it's not so clear that it will work for the DRFPMI, because the signal recycling cavity complicates things.

I need to look at, and meditate over, some Optickle simulations before I say much else about this stuff.

 The idea of introducing a large MICH offset to reduce the PRC finesse might help us to get rid of the transmitted power signal. We might be able to increase enough the line width of the double cavity to make it larger than the ASL length fluctuations. Then we can switch from ASL to the IR demodulated signal without transitioning through the power signal.

  9343   Tue Nov 5 08:44:21 2013 SteveUpdateIOOafter last steering mirror mount swap

 The IOO Angle and IOO Position qpds  were recentered after this entry.

 Suggested corrections in elog entry #9323 are completed:

 1,  last steering mirror mount replaced by Polanski mount

 2,  PSL output shutter mount reconfigured

 IOO qpds are not centered. I failed to connect laptops to 40MARSian network.

Attachment 1: after_Shutter_SMm_installed.png
after_Shutter_SMm_installed.png
Attachment 2: afterlastSMmountswap.jpg
afterlastSMmountswap.jpg
  9342   Tue Nov 5 00:39:43 2013 manasaUpdateIOOIFO alignment tuning

Information acknowledged from Steve:

The last steering mirror mount for IR on the PSL was swapped for a more robust one. Prior to swapping the ibeam positions on the PSL IOO QPDS in ang and pos were recorded.

What I did henceforth:

1. Once the last steering mirror was installed, I walked the beam to restore input pointing using the last 2 steering mirrors. It needed a lot of work in yaw as expected.

2. When the input pointing was recovered, MC locked right away in TEM00. I measured the MC spot positions and compared it with Jenne's measurement made prior to the swap. The spot positions were pretty close.

3. The input pointing was adjusted in pitch and yaw (on the last steering mirror) in small steps. MC alignment was recovered and spot positions were measured each time. After several iterations, the MC spot positions were pretty much centered. I recentered the WFS and reset the WFS offsets. MC is now locked with WFS enabled at ~16900 counts.

MC_spot.png

4. Since the arms were aligned this morning, I used the Y arm as reference and corrected for the input pointing using tip-tilts.

5. Arms locked right away. Note: ASS doesn't seem to be doing it's job. I had to manually align the arms for maximum on TRX and TRY.

6. MICH and PRMI lock were also recovered.

7. I started to check the status with ALS as well. But for reasons unknown, I don't see any ADC counts corresponding to the beat note. Looking at the beatbox there aren't any signs of disconnected cables.  I am saving this as a morning job to fix it.

  9341   Mon Nov 4 23:11:00 2013 JenneUpdateLSCSmall updates to LSC screen

I made some small edits to the LSC screen. 

* When I added columns for the new AS110 PD, I had forgotten to make the Trigger matrix and Power Normalization matrix icons on the screen bigger, so we weren't seeing the last 2 columns in the overview screen.

* I added "show if not zero" oscillator icons to the Sensing Matrix part of the LSC overview screen, so that it's easier at a glance to see that there is an oscillator on.

  9340   Mon Nov 4 18:24:15 2013 KojiUpdateLSCThoughts on Transition to IR

 You have the data. Why don't you just calculate 1/SQRT(TRX)?

...yeah, you can calculate it but of course you don't have no any reference for the true displacement...

  9339   Mon Nov 4 17:08:23 2013 JenneUpdateLSCThoughts on Transition to IR

Gabriele and I talked for a while on Wednesday afternoon about ideas for transitioning to IR control, from ALS. 

I think one of the baseline ideas was to use the sqrt(transmission) as an error signal.  Gabriele pointed out to me that to have a linear signal, really what we need is sqrt( [max transmission] - [current transmission] ), and this requires good knowledge of the maximum transmission that we expect.  However, we can't really measure this max transmission, since we aren't yet able to hold the arms that close to resonance.  If we get this number wrong, the error signal close to the resonance won't be very good.

Gabriele suggested maybe using just the raw transmission signal.  When we're near the half-resonance point, the transmission gives us an approximately linear signal, although it becomes totally non-linear as we get close to resonance.  Using this technique, however, requires lowering the finesse of PRCL by putting in a medium-large MICH offset, so that the PRC is lossy.  This lowering of the PRC finesse prevents the coupled-cavity linewidth of the arm to get too tiny.  Apparently this trick was very handy for Virgo when locking the PRFPMI, but it's not so clear that it will work for the DRFPMI, because the signal recycling cavity complicates things.

I need to look at, and meditate over, some Optickle simulations before I say much else about this stuff.

  9338   Mon Nov 4 15:46:17 2013 JenneUpdateLSCThoughts and Conclusions from last week's PRMI+2arms attempt

5:31pm - This is still a work in progress, but I'm going to submit so that I save my writing so far. I think I'm done writing now.


First, a transcription of some of the notes that I took last Tuesday night, then a few looks at the data, and finally some thoughts on things to investigate.


MICH and PRCL Transfer Functions while arms brought in to resonance (both arms locked to ALS beatnotes):

This is summarized in elog 9317, which I made as we were finishing up Tuesday night.  Here's the full story though.  Note that I didn't save the data for these, I just took notes (and screenshots for the 1st TF).

POP22I was ~140 counts, POP110I was ~100 counts.

MICH gain = -2.0, PRCL gain = 0.070. 

First TF (used as reference for 2-10), PRMI locked on REFL165, Xarm transmission = 0.03, Yarm transmission = 0.05 (both arms off resonance).  MICH UGF~40Hz, PRCL UGF~80Hz.

MICH_40Hz.pngPRCL_80Hz.png

2: X=off-res (xarm not moved), Y=0.13, no change in TF

3: X=off-res (xarm not moved), Y=0.35, no change in TF

4: X=off-res (xarm not moved), Y=0.60, MICH high freq gain went up a little, otherwise no change (no change in either UGF)

5:  X=off-res (xarm not moved), Y=0.95, same as TF#4.

6: X=0.20, Y=1.10 (yarm not moved), same as TF#4

7:  X=0.40, Y=1.30 (yarm not moved), same as TF#4

8:  X=0.70, Y=1.55 (yarm not moved), same as TF#4

9: X=1.40, Y=2.20 (yarm not moved), same as TF#4

10: X=4.0, Y=4.0 (yarm not moved), PRCL UGF is 10Hz higher than TF#4, MICH UGF is 20Hz lower than TF#4.

11: (No TF taken), Xarm and Yarm transmission both around 20!  To get this, MICH FMs that were triggered, are no longer triggered to turn on.  Also, MICH gain was lowered to -0.15 and PRCL gain was increased to 0.1

12: (No TF taken), Xarm and Yarm transmissions both around 40!  The peaks could be higher, but we don't have the QPD ready yet.

After that, we started moving away from resonance, but we didn't take any more transfer functions.


OpLev spectra for different arm resonance values:

We were concerned that the ETMs and ITMs might be moving more, when the arms are resonating high power, due to some optical spring / radiation pressure effects, so I took spectra of oplevs at various arm transmissions.

I titled the first file "no lock", and unfortunately I don't remember what wasn't locked.  I think, however, that nothing at all was locked.  No PRMI, no arm ALS, no nothing.  Anyhow, here's the spectrum:

ALS_noLock.pdf

I have a measurement when the Yarm's transmission was 1, and the Xarm's transmission was 1.75.  This was a PRMI lock, with ALS holding the arms partially on resonance:

ALS_X1pt75Y1.pdf

Next up, I have a measurement when Yarm was 0.8, Xarm was 2.  Again, PRMI with the arms held by ALS:

ALS_X2Y0pt8.pdf

And finally, a measurement when Xarm was 5, Yarm was 4:

ALS_X5Y4.pdf

Just so we have a "real" reference, I have just now taken a set of oplev spectra, with the ITMs, ETMs and PRM restored, but I shut the PSL shutter, so there was no light flashing around pushing on things.  I noticed, when taking this data, that if the PSL shutter was open, so the PRFPMI is flashing (but LSC is off), the PRM oplev looks much like the original "no Lock" spectra, but when I closed the shutter, the oplev looks like the others.  So, perhaps when we're getting to really high powers, the PRM is getting pushed around a bit?

ALS_noLock_noLaser.pdf

Conclusions from OpLev Spectra:  At least up to these resonances (which is, admittedly, not that much), I do not see any difference in the oplev spectra at the different buildup power levels.  What I need to do is make sure to take oplev spectra next time we do the PRMI+2arms test when the arms are resonating a lot. 


Time series while bringing arms into resonance:

PRMI_2arms_29Oct2013_POPrin.png

I had wondered if, since the POP 22 and 110 values looked so shakey, we were increasing the PRCL RIN while we brought the arms into resonance.  You can see in the above time series that that's not true.  The left side of the plot is PRMI locked, arms held out of resonance using ALS.  First the Yarm is brought close to resonance, then the Xarm follows.  The RIN of the arms is maybe increasing a little bit as we get closer to resonance, but not by that much.  But there seems to be no correlation between arm power and RIN of the power recycling cavity.

Alternatively, here is some time series when the arm powers got pretty high:

PRMI_2arms_29Oct2013_POPrin_highArmPowers.png


Possible Saturation of Signals:

One possibility for our locklosses of PRMI is that some signal somewhere is saturating, so here are some plots showing that that's not true for the error and control signals for the PRMI:

PRMI_2arms_29Oct2013_LSCcontrolSignals.png

Here, for the exact same time, is a set of time series for every optic except the SRM.  We can see that none of the signals are saturating, and I don't see any big differences for the ITMs or ETMs in the times that the PRMI is locked with high arm powers (center of the x-axis on the plot) and times that the PRMI is not locked, so we don't have high arm powers (edges of the plot - first half second, and last full second).  You can definitely see that the PRM moves much more when the PRMI is locked though, in both pitch and yaw. 

PRMI_2arms_29Oct2013_OpLevs_highArmPowers.png

DCPD signals at the same time:

PRMI_2arms_29Oct2013_DCPDs.png

NB:  These latest 3 plots were created with the getdata script, with arguments "-s 1067163405 -d 7".  It may be a good idea to take some spectra starting at, say 1067163406, 1 second in, and going for ~2 seconds. (It turns out that this is kind of a pain, and I can't convince DTT to give me a sensible spectrum of very short duration....we'll just need to do this live next time around).


Things to think about and investigate:

Why are we losing lock? 

On paper, is the (will the) optical spring a problem once we get high resonance in the arms? 

Spectra of oplevs when we're resonating high arm power.

What is the coupling between 110MHz and 165MHz on the REFL165 PD?  Do we need a stronger bandpass? 

Why are things so shakey when the arm power builds up?

Why do PRCL and MICH have different UGFs when the arms are controlled by ALS vs. ETMs misaligned?

Does QPD for arm transmissions switching work?  Can we then start using TRX and TRY for control?

What is the meaning of the similar features in both transmission signals, and the power recycling cavity?  Power fluctuation in the PRC due to PRM motion? 

  9337   Mon Nov 4 14:11:23 2013 GabrieleSummaryLSCSimulation of REFL_3f signal when the arms come in

Quote:

Yes, the resonance of the 2nd-order sidebands to the IFO screws up the 3f scheme.

2f (~22MHz) and 10f (~110MHz) are at x 5.6 and x 27.9 FSR from the carrier, so that's not the case.

Could we also see how much gain fluctuation of the 3f signals we would experience when the arm comes into the resonance?

 From the simulation there is no visible change in the gain.

  9336   Mon Nov 4 12:59:43 2013 JenneUpdateIOOMC alignment not so good after PSL output shutter installed

Quote:

  The PSL shutter is reinstalled.

 I'm not sure if Steve bumped something, or if it was just a fluke, but the MC didn't come back very nicely after Steve finished re-installing the shutter.

Earlier today, after Steve locked the PMC, MC trans looked good for over an hour (according to the striptool plot on the wall).  Then, the MC was unlocked for about an hour, presumably while Steve was working, he had the light blocked.  When he finished, the MC transmission was around 5,000 while usually it is around 17,000.  The reflection was around 3.4, rather than a best of below 0.5 (unlocked refl is 4.5).

Using Rana's ezcaservo trick to get the suspensions back to where they were at last good lock usually works (I used to do it by hand though).  However, today, it only got the reflection down to about 2.0.  So, I did the rest of the alignment by hand. 

After I did this, the reflection is down to 0.48.  Engaging the WFS makes the MC much more noisy, so I have them disabled currently. 

I have measured the spots, and if I compare them to the measurements that (I think it was Manasa) took last week, they look pretty bad. 

I think that we need to swap out the 2nd zigzag mirror, and then do a careful MC realignment.  It's certainly not worth doing the work, and then re-doing it after we swap out the zigzag mirror.

MCspots_4Nov2013.png

  9335   Mon Nov 4 12:07:55 2013 DmassOmnistructureGeneralReplaced Broken Drill Bits

I broke a small bit while using the 40m drill press to vent some 1/4-20 screws for the cryo experiment.

I replaced it and refilled the small bit row in the bit index I was using; there were ~10 missing sizes

  9334   Mon Nov 4 11:37:12 2013 SteveUpdateIOOPSL output shutter installed

 

 The PSL shutter is reinstalled.The base plate is delrin for isolation and the mount height is adjustable. The last steering mirror mount to be swapped in is ready. It is sitting on the top of the ITMX optical table cover with SS dogs.

There are two reflected spots on the north side of the Uniblitz shutter. They are coming from the vacuum window. They should be trapped also.

Attachment 1: Shutterisback.jpg
Shutterisback.jpg
  9333   Mon Nov 4 09:03:21 2013 SteveUpdatePSLPMC auto locker

Quote:

 I was working on the electronics bench and what sounded like a huge truck rolled by outside. I didn't notice anything until now, but It looks like something became misaligned when the truck passed by (~6:45-6:50 pm). I can hear a lot of noise coming out of the control room speakers and pretty much all of the IOO plots on the wall have sharp discontinuities.

I haven't been moving around much for the past 2 hours so I don't think it was me, but I thought it was worth noting.

 The PMC auto locker is not set to acquire error message made me lock PMC manually.  Arms  locked instantly: TRY 0.9 V and TRX 0.65 V

Attachment 1: 3dTREND.png
3dTREND.png
  9332   Sun Nov 3 00:05:52 2013 CharlesSummaryISSISS Update - Bout' time

Right near the end of summer, I had an ISS board that was nominally working, but had a few problems I couldn't really sort out. Since I've been back, I've spent a lot of time just replacing parts, trying different circuit topologies and generally attempting to make the board function as I hoped it might in all those design stages. Below is a brief list of some of the problems I've been fixing as well as the first good characterization of the board transfer function that I've been able to get.

We'll start with some of the simple problems and proceed to more complicated ones.

  • The 5V reference I was using to obtain an error signal from some arbitrary DC photodiode readout was only producing ~2.5 V. 
    • Turns out I just need a FET type op-amp for the Sallen-Key Filter that I was using to clean up any noise in the reference output, as the leakage current in a AD829 was causing a significant voltage drop. I put in an OPA140 and everything worked marvelously.
  • The way I set up input grounding (i.e. send a ~0 amplitude signal through the board as an input) passed a few Amps through one of my chips causing it to burn out rather fantastically.
    • There isn't a good way to fix this on the current board (besides just getting rid of the functionality altogether) so my solution so far has just been to redesign that particular sub-system/feature and when we implement the second version of the ISS, the input grounding will be done correctly
  • One of the ICs I'm using, specifically the AD8436 RMS-to-DC converter, causes some super strange oscillations in -5V power line. When this chip is soldered onto the board, the -5V supply jumps between -3V and -10V rather sporadically and the DC power-supply used to provide that -5V says that board is drawing ~600 mA on that particular power line.
    • To date, I don't really have any idea what's going with this chip, and I've tried a lot of things to remedy the problem. My first thought was that I had some sort of short somewhere so I took the chip off the board, cleaned up all the excess solder and flux around the chip's footprint and then meticulously soldered a new chip on (when I say meticulously, it took over an hour to solder 20 little feet. I really really didn't want to short anything accidentally as the chip only comes in a package with ridicously small spacing between the leads). Lo and behold, nothing happened. I still saw the same oscillations in power supply and the board was still drawing between >500 mA on that line. Just to be sure, I soldered on a third chip taking the same amount of care and had the same problems.
    • I went over the schematic in Altium that we used to order the board, and unless the manufacturer made a mistake somewhere, there aren't any incorrectly routed signals would cause, say, two active devices to try setting the voltage of a particular node to different values.
    • I got some QSOP-to-DIP package converters so that I could mess around with the AD8436 on a breadboard to make sure it functioned correctly. I set up an identical circuit to the one on the PCB and didn't see any oscillations in the power supply, both for +-5V and +-15V as the chip can handle both supply voltages. I'm not really sure how to interpret this...
    • I'm still actively trying to figure this particular problem out, but I'm shooting in the dark at this point. 
  • Initial attempts to measure the transfer-function of the board were wrought with failure.
    • I figured out, with Nic's help, that the board needs the 'loop closed' with a significant broadband attenuator (to simulate the plant optics discussed in elog 9331) in order to not have constant railing of the high gain op-amp filter stages. Even after I did this, the measured transfer functions were not at all consistent with simulation. I wasn't sure if it was just a part issue, a design issue or a misunderstanding/bad data collection on my part so I just redesigned the whole servo and stuffed the board with entirely new components from around the 40m. Turns out the newly designed servo behaved more properly, as I will show below.

The above list encompasses all the issues I've had in making the ISS board function correctly. No other major problems exist to my knowledge.

I was able to measure both the open- and closed-loop transfer functions of the servo with the SR785. The results are shown below.

full-op-loop.png

The transfer function with the boosts on caps at a particular value set by op-amp railing, i.e. below 100 Hz, the op-amps are already putting out their max voltage. This is the usual physical limitation when measuring the transfer function of an integrator. We can also see that the measured phase follows the simulated phase above ~300 Hz. The 'phase matching' at low frequency is again do to the op-amp railing in the servo output..

The closed-loop gain is shown below,

full-cl-loop.png

The measured closed-loop gain with the boosts on again matches the LISO simulation quite well except at low frequency where we are limited by op-amp railing. We compare the measured closed-loop transfer function to the desired noise suppression stipulated in my previous elog 9331,

req-vs-meas.png

 And we might hopefully conclude that my servo functions as desired. One should note that the op-amp railing seen in these measurements is not indicative of limitations we might face in some application of the ISS for the following reason. These transfer functions were measured with a 100 mV excitation signal (it is necessary to keep this signal amplitude large enough so that the inherent signal-to-noise ratio of the excitation source is large enough for accurate measurement) which leads to somewhat prompt railing of the op-amps. When the ISS operates to actually stabilize a laser, the input error signal will be much smaller (on the order of a few 10's of mV or less) and will decrease significantly assuming correct operation of the ISS. This means we won't see the same type of gain limitations.

 

What now, you ask?

Aside from the problem with the AD8436 chip, the ISS board seems to be functioning correctly. The transfer functions we have measured are correct to within the component tolerances and all of the various subsystems are behaving as they were designed to. Moving toward the goal of having this system work in situ for the CTN experiment, I need to do the following things,

  • Design a housing for the board -> order said housing and the front panel previously designed
  • Make sure the power supply daughter PCB boards are compatible with the ISS board and can provide power correctly
  • Talk to Evan and Tara about integrating the ISS with their experiment and make sure my board can do everything it needs to in that context.

So close, or so I say all the time 

 

  9331   Sat Nov 2 22:49:44 2013 CharlesUpdateISSCTN ISS Noise Suppression Requirement - Updated 10/27

 Previously in elog 8959, I gave a very simple method for determining the noise suppression behavior of the ISS. Recently, I recalculated this requirement in a more correct fashion and again redesigned the ISS to be used in the CTN experiment.

  • Determining the Requirement

Just as before, the data from PSL elog 1270 is necessary to infer a noise suppression requirement. The data presented there by Evan consists of two noise spectra, 1) the unstabilized RIN presently observed in the CTN experiment readout and 2) the theoretical brownian noise produced by thermal processes in the mirror coating+substrate. The statement "TF_mag = (Unstabilized RIN) / (Calculated Brownian Noise Limit)", where TF_mag refers to the required open-loop gain of the ISS, is actually a first order approximation of the 'required' noise suppression. In fact if we wanted the laser noise to be suppressed below the calculated brownian noise level, it is more correct to say 

        Closed-loop ISS gain = (Calculated Brownian Noise Limit) / (Unstabilized RIN)

As this essentially gives a noise suppression spectrum i.e. a closed-loop gain in linear control theory. Below is a very simple block diagram showing how the ISS fits into the CTN experiment. The F(f) block represents my full servo board.

    ISS_path.png

Some of the relevant quantities involved:

            plant-quant_1.png

            plant-quant_2.png

So looking at the block diagram, our full closed-loop transfer function is given by,

cl-loop.png

So then to determine the required F(f), i.e. the required transfer function for my servo, we consider the expression 

               requirement.png

The plant transfer function is simply Plant = (C(f) * a * P * A) ~ 0.014 V/V, where I have ignored the cavity pole around 97 kHz as our open-loop transfer function ends up crossing unity gain around 10 kHz. In the above, I have included what I call a 'safety factor' of 10. Essentially, I want to design my servo such that it suppresses noise well beyond what is actually required so that we can be sure noise contributions to experiment readouts are not significantly influenced by the laser intensity noise.

  • Proposed Servo Design

Using the data Evan reported for the brownian noise and free-running RIN, I came up with an F(f) to the meet the requirement as shown below.

CTN_TF_req-vs-proposed.png

 Where the blue curve includes the safety factor mentioned before. This plot just demonstrates that using my modular ISS design, I can meet the given noise suppression requirements.

To be complete, I'll say a little more about the final design.  As usual, the servo consists of three stages. The first is the usual LP filter that is always 'on' when the ISS loop is closed. The boosts I have chosen to use consist of an integrator with a single zero and a filter that looks somewhat like a de-whitening filter. The simulated open-loop transfer functions are shown below.

switching-filters.png

 

 

 

 

 

 

 

 

  9330   Sat Nov 2 19:36:15 2013 CharlesUpdateGeneralPossible misalignment?

 I was working on the electronics bench and what sounded like a huge truck rolled by outside. I didn't notice anything until now, but It looks like something became misaligned when the truck passed by (~6:45-6:50 pm). I can hear a lot of noise coming out of the control room speakers and pretty much all of the IOO plots on the wall have sharp discontinuities.

I haven't been moving around much for the past 2 hours so I don't think it was me, but I thought it was worth noting.

  9329   Fri Nov 1 19:09:01 2013 rana, evanConfigurationPSLPMC reflected beam nonsense

 While looking at the PMC REFL beam for the AOM diffracted beam, we noticed that although only one beam exists between the PMC and the first steering mirror, there are two afterwards and they both go to the PMC REFL  RFPD!!! This is madness. We only want one beam on our PDH diode.

The reason that we have two beams is that that first steering mirrors is actually a (W1-PW-1025-UV-1064-45P) non-wedged window with an AR coating on only one side. So two beams come out of it. There is a terrible and floppy and illegal anodized aluminum dump close to this beam which *someone* probably intended to use as a "scraper" to get rid of one of the beams.

Black anodized aluminum is a horrible beam dump material at 1064 - its about as grey as Steve's chair. And its so soft that it scatters light back into the PMC and makes more acoustic noise. And it is mounted so poorly (only one screw) that it can easily be bumped and twist and miss the beam. Punchline: only use anodized aluminum dumps for stray light around cameras or for HeNe for OL. Its NOT allowed anywhere where we care about interferometry of NIR beams.

It was also set to dump the dimmer beam. On Monday, we should order ~5 W1 and get them with a wedge of 1-2 deg. Then we use a black glass dump for the dim beam and orient the bright one to hit the REFL camera and the PMC REFL PD.

For the weekend, I have adjusted the crappy grey aluminum flapper to catch the bright beam so that the PMC REFL image no longer shows the interference fringe of two beams. Lets see how the PMC drifts over the next 3 days.

  9328   Fri Nov 1 18:59:41 2013 EvanConfigurationISSAOM cabling

[Rana, Nic, Evan]

We did some work today on getting the AOM back up and running so that we can implement an SR560-based ISS.

We've removed the 18 AWG wire that was previously used to power the driver and have replaced it with a 12 AWG twisted pair (black and white, enclosed in a single gray cladding). This pair runs into the PSL rack's 24 V terminal block with a 2 A fuse. We've also replaced the cable connecting the AOM to the driver; it's now RG405.

Also disconnected the power to the old Kalmus FSS crystal driver box and turned it off. It was powered illegally. Also disconnected the power connection between the Sorensen and the old ISS AA chassis since it was wired directly without any fuse (which is a code violation). It will stay off until someone uses a proper fuse and wiring to hook it back up.

Attachment 1: aom.jpg
aom.jpg
Attachment 2: aom_driver.jpg
aom_driver.jpg
Attachment 3: aom_driver_power.jpg
aom_driver_power.jpg
Attachment 4: 20131101_170120.jpg
20131101_170120.jpg
  9327   Fri Nov 1 17:44:06 2013 KojiSummaryLSCSimulation of REFL_3f signal when the arms come in

Yes, the resonance of the 2nd-order sidebands to the IFO screws up the 3f scheme.

2f (~22MHz) and 10f (~110MHz) are at x 5.6 and x 27.9 FSR from the carrier, so that's not the case.

Could we also see how much gain fluctuation of the 3f signals we would experience when the arm comes into the resonance?

  9326   Fri Nov 1 17:01:46 2013 GabrieleSummaryLSCSimulation of REFL_3f signal when the arms come in

 I simulated how the 3f signal is affected by the resonance condition of the arms.

To keep it simple, I only simulated a double cavity. The attached plot shows the result. In x there is the arm cavity detuning from resonance (in log scale to show what happens close to the 0 value). In the y axis there is the PRC detuning. So every vertical slice of the upper plot gives a PDH signal for a given arm detuning. The bottom plot shows the power build up inside the arm, which is dominated by the carrier.

refl_3f_I_vs_cavity_tuning_noresf2.png

The 3f signal is not perturbed in any significant way by the arm resonance condition. This is good and what we expected.

However, in this simulation I had to ensure that the 1f sidebands are not perfectly anti-resonant inside the arms. They are indeed quite far away from resonance. If the modulation frequency is chosen in order to make the 1f sidebands exactly ant-resonant, the 2f will be resonant. This screws up the signal: REFL_3f is made of two contributions of equal amplitude, one on the PRC sidebands resonance and the other on the PRC carrier resonance. When the arm tuning goes to zero, these two cancels out and there is no more PDH...

refl_3f_I_vs_cavity_tuning.png

However, this is a limit case, since the frequency show match perfectly. If the modulation frequency is few arm line widths away from perfect anti-resonance, we have no problem.

  9325   Fri Nov 1 09:45:32 2013 SteveUpdateIOObeam dumps to be find

Quote:

Quote:

Steve, can you please put 2 mounted  razor dumps near the WFS for this purpose??    

            Tuesday: Razor dumps are waiting for you.

 I couldn't find any dumps near the WFS. Koji looked. I looked twice. Maybe they are spooky and absorbing all of the light?

The MC alignment was bad and the WFS were making it drift. Koji aligned the beam into the PMC. I then restored the MC suspensions to where they were 8 days ago (back when the transmission and reflection were good). With the WFS OFF, this gave us a MC trans ~ 16000. With WFS ON it goes to 17500 which is about as good as its been over the last 80 days.

I centered the beam on the WFS with the MC unlocked and also centered the beam on the whole WFS path (it was near clipping between WFS 1 & 2). Also for some reason that beamsplitter which steers the beam onto WFS1 is a R=33% (!? why is this not a R=50% ??).

Steve, please swap this out to a BS1-1064-50-1025-45S if we have one sitting around. If not, we want to add this to the CVI purchase list, but not buy until we get a bigger list together.

I also centered this newly aligned beam into the IMC onto the PSL QPDs. We should now use these as a pointing reference for the beam into the IMC.

While doing this I noticed that the beam was almost clipping on the Uniblitz shutter used to block the PSL beam. That shutter is mounted too short and was also not centered horizontally. I removed it for now so that Steve can find a more adjustable mount for it and put it back into play. The beam going into the IMC is BIG, so you have to very careful when centering the shutter. Might be that we cannot leave it at 45 deg and still get a big enough aperture.

Note #3 for Steve: please also replace the mount for last steering mirror into the IMC with a Polanski or a Superman, that black Ultima is no good. Also the dogs must be steel - no aluminum dogs for our sensitive places.

No wonder they could not find the beam dumps. Last night was Haloween. They should of just said: Trick or treat! where are the beam dumps?

Attachment 1: razor_beam_dumps.jpg
razor_beam_dumps.jpg
  9324   Thu Oct 31 21:22:00 2013 rana, kojiSummaryIOOmodulation beat note in MC servo

I hooked up the 4395 to the MC servo board test out (TP2A) and looked at the spectrum using our new SPAG4395.py script. We noticed a huge peak at ~3.8 MHz and correctly guessed that it was due to the beat between the MC modulation frequency 29.5 MHz and 3*f1 (~33 MHz).

So we tuned the Marconi for the main mod. from 11065910 to 11066099 Hz and saw the beat note disappear (to within the 1 Hz tuning precision of our Marconi).

New MC length tuning method! Alert the LA Times!

13031.png13031_200.png13031_200b.png

My conjecture is that this temperature dependent mismatch between the modulation frequency (f1) and the MC length  is what leads sometimes to our nasty saturating PC DRIVE signal. TBD.

  9323   Thu Oct 31 20:05:48 2013 RANAUpdateIOOMode Cleaner Tune-UP

Quote:

Steve, can you please put 2 mounted  razor dumps near the WFS for this purpose??    

            Tuesday: Razor dumps are waiting for you.

 I couldn't find any dumps near the WFS. Koji looked. I looked twice. Maybe they are spooky and absorbing all of the light?

The MC alignment was bad and the WFS were making it drift. Koji aligned the beam into the PMC. I then restored the MC suspensions to where they were 8 days ago (back when the transmission and reflection were good). With the WFS OFF, this gave us a MC trans ~ 16000. With WFS ON it goes to 17500 which is about as good as its been over the last 80 days.

I centered the beam on the WFS with the MC unlocked and also centered the beam on the whole WFS path (it was near clipping between WFS 1 & 2). Also for some reason that beamsplitter which steers the beam onto WFS1 is a R=33% (!? why is this not a R=50% ??).

Steve, please swap this out to a BS1-1064-50-1025-45S if we have one sitting around. If not, we want to add this to the CVI purchase list, but not buy until we get a bigger list together.

I also centered this newly aligned beam into the IMC onto the PSL QPDs. We should now use these as a pointing reference for the beam into the IMC.

While doing this I noticed that the beam was almost clipping on the Uniblitz shutter used to block the PSL beam. That shutter is mounted too short and was also not centered horizontally. I removed it for now so that Steve can find a more adjustable mount for it and put it back into play. The beam going into the IMC is BIG, so you have to very careful when centering the shutter. Might be that we cannot leave it at 45 deg and still get a big enough aperture.

Note #3 for Steve: please also replace the mount for last steering mirror into the IMC with a Polanski or a Superman, that black Ultima is no good. Also the dogs must be steel - no aluminum dogs for our sensitive places.

Attachment 1: drifty.png
drifty.png
  9322   Thu Oct 31 15:34:28 2013 manasaUpdateIOOMC not happy since yesterday

Quote:

 

8 day minute trend of some of the IMC alignment signals.

That step ~2 days ago in the WFS2 yaw control signal shows that I didn't do such a good job on yaw.

Nic is going to come over some time and give us a new Gouy telescope that let's us have bigger beams on the WFS. At LLO, Hartmut demonstrated recently how bigger beams can reduce offsets somehow...mechanism TBD.

Also, we must angle the WFS and figure out how to dump the reflections at the same time that we rework the table for the telescope.

Steve, can you please put 2 mounted  razor dumps near the WFS for this purpose??    

            Tuesday: Razor dumps are waiting for you.

 

The MC has not been able to hold lock for over a couple of hours since yesterday. I aligned the MC yesterday (elog 9320) and it lost lock in a couple of hours. I realigned the MC again around noon today only to see it drifting and lose lock again.

I have attached the MC trend for the 2 hours when the MC drifted slowly from its happy to sad state.

 

 

Attachment 1: MC_drift.png
MC_drift.png
  9321   Thu Oct 31 00:06:45 2013 masayukiUpdateComputer Scripts / Programscds fft,tf,offsets

I wrote the scripts for cdsutils.

1. fft

usage: fft.py [-h] [-c N_CYCL] [-a N_AVG] freq channel [channel ...]

Measures the frequency content of one or more NDS channels
at the specified frequecy.
    The measurement results are  magnitude, phase, real part imaginary part, and the standard deviation of the real and imaginary parts.

positional arguments:
  freq        measurement frequency
  channel     list of measurement channel

optional arguments:
  -h, --help  show this help message and exit
  -c N_CYCL   define number of cycles. Default is 10
  -a N_AVG    define number of average. Default is 100

2. tf

usage: tf.py [-h] [-c N_CYCL] [-a N_AVG] [--dB] freq input output

Measures the transfer funtion of one NDS channels pair  at the specified frequecy.
    The measurement results are the coherence, magnitude, phase, real part, imaginary part, and the standard deviation of the real and imaginary parts.

positional arguments:
  freq        measurement frequency
  input       list of measurement input channel
  output      list of measurement output channel

optional arguments:
  -h, --help  show this help message and exit
  -c N_CYCL   define number of cycles. Default is 10
  -a N_AVG    define number of average. Default is 100
  --dB        print the amplitude in dB

 

3.offsets

usage: offset.py [-h] [-t SECONDS] filterbank [filterbank ...]

Zero the offset of one or more filterbank. Take average of OUT16 channel, and put (-1 * average / gain) into offset.  After change offsets, it will turn on the offset.
    example usage) offset -d 25 C1:LSC-AS55_Q C1:LSC-AS55_I

positional arguments:
  filterbank  list of filterbank to zero offset

optional arguments:
  -h, --help  show this help message and exit
  -t SECONDS  define the averaging time. default value is 10

I put these scripts in /opt/rtcds/cdsutils/trunk/lib/cdsutils.
I couldn't put them into cds library, but I will left tomorrow to Japan...

So please help me Jamie or Joe !!

 

 

By the way,

I modified the LSCoffset script  (script)/LSC/LSCoffsets.py.

usage: LSCoffsets.py [-h] [-d SECONDS] [--PDlist]

 Zero the offsets of LSC PDs. During taking average, we will close the PSL and green shutter. After zeroing, it will turn on the offsets switch.
    example usgae) LSCoffset2.py -d 20    

optional arguments:
  -h, --help  show this help message and exit
  -d SECONDS  define the averaging time. default value is 10
  --PDlist    print PD list for LSC and exit

i made new directory 'offset_backup' for old offset script. I move all old offset script into there.

But unfortunately that cannot use right now, because cds offsets script is not available yet...

  9320   Wed Oct 30 16:46:17 2013 manasaUpdateIOOMC aligned

MC has not been very happy since last night. 

What I did to fix this:

1. Disabled autolocker and WFS and aligned the MC to bring MC REFL down to <0.50

2. When I re-enabled autolocker, MC was losing lock everytime WFS turned ON.

3. I relocked MC, measured the spot positions and moved MC spot positions by running /opt/rtcds/caltech/c1/scripts/ASS/MC/mcassMCdecenter 
and 
/opt/rtcds/caltech/c1/scripts/MC/moveMC2/

4. I reset the WFS offsets by running /opt/rtcds/caltech/c1/scripts/MC/WFS/WFS_FilterBank_offsets

5. MC is locked and looks happy right now with REFL DCMON at ~0.5

 

  9319   Wed Oct 30 14:58:44 2013 SteveHowToGeneralshutting down a computer

 

 We have tried to ssh into c1iscey yesterday morning. It just did not work. We have just tried it again (now) and it did work.

Lesson learned: always shut down the computer from a TERMINAL Do not turn it off by the manual power switch.

  9318   Wed Oct 30 08:32:55 2013 EvanHowToComputer Scripts / Programsmatplotlibrc

FYI, you can trick out matplotlib by creating a matplotlibrc config file. This allows you to set defaults for plot size, trace color, fonts, grids, etc., analogous to what is achieved in ATF:1840 for Matlab.

Note also that matplotlib supports LaTeX by default (if you have LaTeX installed), which means, for example, that you can include true square roots on your spectral densities:

plt.ylabel('Voltage spectral density (V/$\\sqrt{\\mathrm{Hz}}$)')

Since the backslash is used for escape characters in python, you must escape LaTeX backslashes.

For maximum effect, you can set the following lines in your matplotlibrc file:

text.usetex = True

text.latex.preamble = \usepackage{txfonts}

Then all text and mathematics in your plot will be sent through LaTeX for processing and will appear in Times.

Also, why is the conversion from watts to volts V = 50 * sqrt(W) and not V = sqrt(50 * W)?

  9317   Wed Oct 30 03:36:51 2013 JenneUpdateLSCMICH and PRCL UGFs change with ALS enabled

Masayuki was able to hold both arms off-resonance with ALS long enough for me to lock the PRMI (arms still held off resonance), and take a set of transfer functions.

MICH gain is still -2.0, PRCL gain is still 0.070, which, with the ETMs misaligned, gave me UGFs of 70 for MICH and 180 for PRCL.

Now, however, with the ETMs aligned, but arms held off resonance with ALS, the UGFs have been lowered by a factor of 2 in frequency!  What is doing this??  MICH is now 40 Hz, and PRCL is now 80 Hz.

We measured the MICH and PRCL loops for several arm powers, and there was no change, at least until the arms were both resonating with powers of ~4 . 

After misaligning the ETMs, I remeasured the loops, and the UGFs went back up to where they started.

  9316   Wed Oct 30 03:33:17 2013 RanaUpdateLSCLSC demod boards need some thought

 

 0309.png

I worked on the script SPAG4395A.py tonight with Masayuki's help. This sets up the parameters on the Agilent 4395A and then acquires the spectrum data. It had a couple of bugs before: no matter what channel you requested, you always got channel R. It also would disobey any requests to reduce the attenuation and left the Auto Atten ON. The version now in the SVN allows you to choose the channel and the attenuation.

It then makes this plot using matplotlib. The attached image is from the REFL165 pickoff at a time tonight when the arm powers were ~5-10. I have converted the spectrum from RF electrical Watts into Volts (V = 50*sqrt(W)). To go from the analyzer input to the demod board input we should scale this spectrum by a factor of ~15 (to account for the 20 dB from the coupler and the 3 dB of the splitter and a little more for losses). On the oscilloscope we see Vpp ~5 mV, so that's ~75 mVpp at the output of the BBPD which we're using for REFL165. Perhaps we can handle another factor of ~2-3 ? I'm not sure what we have in terms of linearity measurements on this thing.

EDIT: Evan is right, its V = sqrt(50*W), not V = 50*sqrt(W). ignore y-axis above

  9315   Wed Oct 30 01:53:52 2013 JenneUpdateIOOMode Cleaner relocked

The MC (mostly MC2) decided a few minutes ago to move, so I put the SUSPIT and SUSYAW numbers back where they were, and the tweaked up the alignment from there to get a low MC REFL DC number.  Now the MC is staying locked again, after 20 minutes of not.

  9314   Wed Oct 30 01:44:13 2013 JenneUpdateLSCMICH and PRCL gains adjusted (Config file saved)

Quote:

I am now taking transfer functions of the MICH and PRCL loops to make sure that we have the gains about right.

 I have set the PRCL UGF to be about 180Hz, and the MICH UGF to be about 70 Hz. 

This is with locking on REFL165 I&Q, with MICH gain of -2.0 and PRCL gain of 0.70 . 

The PRCL loop only has about 30 degrees of phase margin, and is not near the top of its phase bubble.  During the day, I need to look at why we don't have more phase near 200 Hz.

  9313   Wed Oct 30 01:22:56 2013 JenneUpdateLSCREFL 165 demod phase adjusted

Quote:

For tonight, Rana has put a coax 100 MHz highpass filter on the input to the REFL165 demod board.

 This of course changes our demod phase.  Rana plotted a 4th order elliptic filter in Matlab, and from the plot determined that we should expect around 60 degrees of difference in our phase. 

To actually set the phase, I locked PRMI on AS55Q and REFL33I (MICH gain = -8.0, PRCL gain = +0.05, with 1's in the matrix elements).  I then turned on the PRCL oscillation notch (564 Hz), and turned on the sensing matrix's drive at that frequency, and looked at the spectrum of REFL165. 

The previous REFL165 demod phase was 96 degrees, so I was looking around either 36 degrees or 156 degrees.  The phase that minimized the peak in the Q signal while driving PRCL was 37.5 degrees.  Good work Matlab/Rana.

I then looked at the transfer functions between REFL33 and AS55 and REFL165, to see if there were any sign flips that happened.  There were not.  As expected, it was just a little extra phase delay.

I was able to lock PRMI with REFL 165 again after this phasing, and I am now taking transfer functions of the MICH and PRCL loops to make sure that we have the gains about right.

  9312   Wed Oct 30 00:02:25 2013 JenneUpdateLSCLSC demod boards need some thought

As we are meditating on things to look at for PRMI + 2 arms, Rana brought up the question of the demod board situation. 

We then found this table on the wiki (LSC demod boards) that indicates that all of the demod boards were originally given lowpass filters, no matter the demodulation frequency.  Back in September, I switched out the low pass filter for a bandpass filter in POP110, and put in the same bandpass when putting together AS110 (elog 9100).  So, the 11MHz diodes are probably okay with lowpasses, and the 110 diodes are okay, but we need to think about all the other ones. 

We should probably do a first guess by putting in a bandpass filter, but then simulate and measure to figure out what our requirements are for attenuation at the non-demodulation frequencies for each board.

The SXBPs from Minicircuits look pretty good, but there are lots of options on their website.

 

For tonight, Rana has put a coax 100 MHz highpass filter on the input to the REFL165 demod board.

  9311   Tue Oct 29 22:33:57 2013 ranaUpdateSUSETMY sensors compared to ETMX

Quote:

I've stopped the process of c1tst again to make it get better. At 9:20, I also went and opened the front rack door (the back one was already open). One reason its hot may be that the exhaust vents on the top of c1iscey are blocked by one of the custom multi-pin adaptor boxes. In the morning, we should drop the computer down by 1 or 2 notches in the rack so that it can air cool itself better. Make sure to poweroff the computer from the terminal before moving it though.

 After some torture Masayuki admitted that he and Steve ignored this elog and just turned off the power button. He blames Steve entirely.

to keep from damaging our computers and our data, NEVER DO THAT.

Scratch-n-SniffStinkyDooBabyShowerGame1.jpg

  9310   Tue Oct 29 18:54:36 2013 JamieConfigurationComputer Scripts / Programsfixing python-matplotlib from LSCSOFT

Quote:
controls@pianosa:~ 0$ cat /etc/apt/preferences.d/pin_python-matplotlib
Package: python-matplotlib
Pin: release a=lucid
Pin-Priority: 1000 

I forgot that there were a couple of different matplotlib packages that all needed to be pinned.  To be safe I decided to just pin all packages to the lucid versions.  This will still allow us to install lscsoft packages that are not ubuntu, but it will always prefer packages from lucid instead.  Here's the new pinning file:

controls@pianosa:~ 0$ cat /etc/apt/preferences.d/pinning 
Package: *
Pin: release a=lucid
Pin-Priority: 1000
controls@pianosa:~ 0$ 

  9309   Tue Oct 29 18:14:52 2013 JamieConfigurationComputer Scripts / Programsfixing python-matplotlib from LSCSOFT

Jenne just discovered an issue with the python-matplotlib package that I knew was coming but forgot about.

We pull packages from the LSCSOFT Debian "squeeze" archive, which is a convenient way for us to install LIGO data analysis software.  There are no LSCSOFT archives for Ubuntu, and Debian "squeeze" is the closest supported distribution to Ubuntu 10.04 "lucid", which is what we are using.

DASWG recently added python-matplotlib to the LSCSOFT squeeze archive.  The version they added (1.0.1-3) supersedes the version in lucid, so by default apt wants to install it.  However, the LSCSOFT version is compiled against a newer version of some standard libraries, so it won't function on our system and seg faults.

The solution (a solution) is to use apt "pinning" to pin the package to the lucid version that works.  I've added the following file on all the 10.04 workstations to prevent the package from upgrading to the LSCSOFT version:

controls@pianosa:~ 0$ cat /etc/apt/preferences.d/pin_python-matplotlib
Package: python-matplotlib
Pin: release a=lucid
Pin-Priority: 1000

 

  9308   Tue Oct 29 16:51:31 2013 JenneUpdateCDSLSC test points were used up

Masayuki was concerned that some LSC channels were giving him all zeros.  After seeing the error in the terminal window running dataviewer (it said something like 'daqd overloaded'), I checked the lsc model, and sure enough, all the test points were used.

So, I found an entry by Jamie (elog 8431) where he reminds us how to clear the test points.  I followed the instructions, and now we're seeing real data (not digital zeros) again.

  9307   Tue Oct 29 10:51:16 2013 MasayukiUpdateSUSETMY sensors compared to ETMX

[Steve, Masayuki]

We lowered the c1iscey machine to make space upside of the computer for heat flow. 

First we turned off the computer. And then we droped the computer down by 1  notches in the rack. Now the upside and downside spaces are almost same. We restarted the computer after that and we leave the door open.

 


  9306   Mon Oct 28 21:33:55 2013 RANAUpdateIOOMode Cleaner Tune-UP

 

8 day minute trend of some of the IMC alignment signals.

That step ~2 days ago in the WFS2 yaw control signal shows that I didn't do such a good job on yaw.

Nic is going to come over some time and give us a new Gouy telescope that let's us have bigger beams on the WFS. At LLO, Hartmut demonstrated recently how bigger beams can reduce offsets somehow...mechanism TBD.

Also, we must angle the WFS and figure out how to dump the reflections at the same time that we rework the table for the telescope.

Steve, can you please put 2 mounted  razor dumps near the WFS for this purpose??    

            Tuesday: Razor dumps are waiting for you.

 

Attachment 1: Untitled.png
Untitled.png
  9305   Mon Oct 28 18:57:27 2013 MasayukiHowToLSCread 'scope and spectrum analyser datas

 

The command to get the data from spectrum analyzer right now

From command line, put ./netgpibdata -i 192.168.113.108 -d AG4395A -a 17 -f meas01

(EDIT JCD:  You must first be in the correct folder:  /opt/rtcds/caltech/c1/scripts/general/netgpibdata/)

(EDIT JCD again: "meas01" in the command line instruction will be the name of the filename.  Also, the output file meas01.dat has a comment in the first line that must be deleted before you can plot the data.  This sucks, and we should write a script to strip that line, then make nice plots.)

Please take notice that although IP address of AG4395A is same as written in the help of netgpibdata, the GPIB address is not same. It's 17.

 

How to use  'scope from control room.

Open the browser. Put the IP adress of 'scope (192.168.113.25) into adrress bar of the browser. If it's on the network, below screen will open.

You can control 'scope, get the data, and so on from control room.

Please take notice that Google Chrome cannot connect the 'scope. So you have to use the Firefox or other browser.

  9304   Mon Oct 28 14:24:01 2013 MasayukiUpdateLSC'scope and spectrum analyser for REFL165

 

 I connected the 'scope between REFL165 output and demod board input. I split the signal from coupler using the splitter (Mini-Circuits ZFSC-2-5). One signal is going to 'scope CH1 and the other is going to spectrum analyzer. I connected the 'scope to 40MARS. The IP adress is 192.168.113.25. I connected that by cabling from 1X2.

 

  9303   Mon Oct 28 14:12:48 2013 manasaUpdateIOOMode Cleaner relocked

Quote:

 The MC had been unlocked for the last 4 hours and was crying out to me so I gave it some attention. Its happier now.

From the trend (AtM #1), I saw that the MC2 suspension has moved by ~10 microradians. Since the MC cavity divergence angle is lambda/(pi*w0) ~ 200 microradians, this isn't so much, but enough to cause it to lock on bad modes sometimes. Attackmint too shows that there's not much in monotonic drift over the last 40 nights.

I moved back MC2 to its old alignment with these commands:

ezcaservo -r C1:SUS-MC2_SUSPIT_INMON -s -1017 -g 0.0009 C1:SUS-MC2_PIT_COMM -t 300

ezcaservo -r C1:SUS-MC2_SUSYAW_INMON -s 490 -g 0.0009 C1:SUS-MC2_YAW_COMM -t 332

Then I went out to the table and aligned the beam into MC using the last two steering mirrors good enough so that the WFS coming on doesn't make the visibility any better. In this nominal state, I unlocked the MC and then aligned the reflected beam onto the center of the LSC PD as well as the WFS. The beam on the first WFS is a little small - next time someone wants to improve our Gouy phase telescope, we might try to make it bigger there. On the LSC PD, the beam was off-center by a few hundred microns.

Masayuki was running LAN cables near the MC2 chamber. This caused the MC2 suspension to move and unlocked the MC. I looked at the long term (2 days) and short term (2 hours) trend of the MC suspensions. I restored the old alignment as described above using ezcaservo.

C1:SUS-MC2_SUSPIT_INMON was restored to 1020 and C1:SUS-MC2_SUSYAW_INMON was restored to 490.

Attachment: Dataviewer trend (2 hours)

Attachment 1: Screenshot-Untitled_Window-3.png
Screenshot-Untitled_Window-3.png
ELOG V3.1.3-